# GrokFormer: Graph Fourier Kolmogorov-Arnold Transformers

### v2

@@ Abstract
Graph Transformers (GTs) have demonstrated remarkable performance in graph representation learning over popular graph neural networks (GNNs). However, self–attention, the core module of GTs, preserves only low-frequency signals in graph features, leading to ineffectiveness in capturing other important signals like high-frequency ones. Some recent GT models help alleviate this issue, but their flexibility and expressiveness are still limited since the filters they learn are fixed on predefined graph spectrum or order. To tackle this challenge, we propose a Graph Fourier Kolmogorov-Arnold Transformer (GrokFormer), a novel GT model that learns highly expressive spectral filters with adaptive graph spectrum and order through a Fourier series modeling over learnable activation functions. We demonstrate theoretically and empirically that the proposed GrokFormer filter offers better expressiveness than other spectral methods. Comprehensive experiments on 10 real-world node classification datasets across various domains, scales, and graph properties, as well as 5 graph classification datasets, show that GrokFormer outperforms state-of-the-art GTs and GNNs. Our code is available at https://github.com/GGA23/GrokFormer.
@@ 1 Introduction
Graph neural networks (GNNs), which jointly encode graph structures and node features, have been emerging as an effective generic tool for graph-structured learning problems. Despite their effectiveness, popular GNNs are often limited by issues like over-smoothing (Oono & Suzuki, 2020) and over-squashing (Topping et al., 2022). On the other hand, graph Transformers (GTs) use a Transformer-based architecture (Vaswani et al., 2017) to learn graph representations. Due to its strong capability in capturing long-range dependencies among graph nodes, it offers a potential solution to address the issues in popular GNNs.
@@
To address this issue, inspired by polynomial GNNs (e.g., ChebyNet (Defferrard et al., 2016), GPRGNN(Chien et al., 2021), BernNet(He et al., 2021) and JacobiConv (Wang & Zhang, 2022)), some recent methods are dedicated to capturing various frequency signals by order-$K$ polynomial approximation. For example, FeTA (Bastos et al., 2022) and PolyFormer (Ma et al., 2024) learn the coefficients for order-$K$ polynomial bases (e.g., Chebyshev, Monomial, or Bernstein basis) through the self-attention mechanism. However, these polynomial filters are typically approximated via $K$ predefined bases with specific frequency responses as illustrated in Figure LABEL:fig:movtivation(a), and thus, they have a receptive field of size $K$ in information passing, leading to a locality modeling and limited flexibility and expressivity. Consequently, they are often unable to fit complicated graph filters well, as shown by the failure of the prevalent Bernstein polynomial in fitting a low-comb filter in Figure LABEL:fig:movtivation(b). To effectively encode such spectrum information and achieve a more global graph modeling, Specformer (Bo et al., 2023) performs self-attention over the $N$ eigenvalues after positional encoding to build learnable filter bases. As shown in Figure LABEL:fig:movtivation(a), the frequency response of its filter bases can be arbitrary on the first-order graph Laplacian spectrum. Although it shows impressive effectiveness, its filter learning has a computational complexity of $O(N^{2})$, making it difficult to simultaneously capture higher-order spectral information, and thus misses some important frequency components embedded in higher-order spectrum. Therefore, the learning capacity of the Specformer filter is limited to the specific first-order spectrum and thus it struggles to fit the complicated low-comb filter in Figure LABEL:fig:movtivation(b). Accordingly, the key question we ask here is: can we have a GT that efficiently and flexibly extracts rich frequency signals across the multi-order spectrum of the graph Laplacian?
@@
To answer this question, we propose a novel GT model, called Graph Fourier Kolmogorov-Arnold Transformers (GrokFormer), which provides an efficient approach for learning order- and spectrum-adaptive graph filter for GTs. In particular, motivated by Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024), GrokFormer leverages learnable activation functions modeled as Fourier series over $K$-order graph Laplacian spectrum, producing $K$ adaptive filter bases. As shown in Figure LABEL:fig:movtivation(a), these bases can capture any frequency response across the spectrum of both low and high orders (i.e., from 1st to $K$th orders). Furthermore, we devise learnable order coefficients to assign varying importance to the $K$ filter bases, enabling an adaptive adjustment in fitting the graph order. The resulting filter in GrokFormer is adaptive/learnable in both graph spectrum and order, having better adaptivity than existing popular filters, as shown in Table 1. In doing so, GrokFormer filter can more flexibly capture a broader range frequency responses, offering significantly better expressiveness than other filters (see Figure LABEL:fig:movtivation(b)). The main contributions are as follows:
@@
  • We propose a novel GT model, GrokFormer, that can effectively capture a wide range of frequency signals in an order- and spectrum-adaptive manner. To the best of our knowledge, this is the first GT model that has a learnable filter in both graph spectrum and order.
@@ 2 Related Work
Graph Neural Networks. Existing GNNs are mainly divided into two main streams: spatial-based and spectral-based methods. Spatial-based GNNs, like GCN (Kipf & Welling, 2017) and GAT (Veličković et al., 2018), update node representations by aggregating information from neighbors. By stacking multiple layers, they may learn long-range dependencies but suffer from over-smoothing and over-squashing. Some improved spatial methods, such as H2GCN (Zhu et al., 2020) and HopGNN (Chen et al., 2023c), propose to combine first-order and higher-order neighborhood representations. Other studies (Xu et al., 2019; Dong et al., 2021) point out from a spectral perspective that GCN only considers the first-order Chebyshev polynomial, which acts as a low-pass filter. Subsequently, various spectral-based GNNs have been proposed, such as, GPRGNN (Chien et al., 2021), BernNet (He et al., 2021), and JacobiConv (Wang & Zhang, 2022) learn arbitrary graph spectral filters by order-$K$ polynomial approximation. HiGCN (Huang et al., 2024) uses Flower-Petals Laplacians in simplicial complexes to learn polynomial filters across varying topological scales. However, the information passing in these polynomial models is local, and their filters with fixed bases have limited learning ability.
Graph Transformers. Compared to GNNs, the attention weights in Transformers can be viewed as a weighted adjacency matrix of a fully connected graph, capturing long-range dependencies. Some GTs combine both and are popular in graph representation learning, such as SAN (Kreuzer et al., 2021), Graphormer (Ying et al., 2021), GPS (Rampášek et al., 2022), SAT (Chen et al., 2022), NodeFormer (Wu et al., 2022), NAGphormer (Chen et al., 2023a) and SGFormer (Wu et al., 2024) are proposed by incorporating various graph structural information into the Transformer architecture. However, these GTs are limited by the inherent low-pass nature of the self-attention mechanism (Bastos et al., 2022). Advanced GTs have increasingly focused on capturing various frequency signals to tackle the issue. SignGT (Chen et al., 2023b) designs a signed self-attention mechanism to capture low- and high-frequency signals. FeTA (Bastos et al., 2022) and PolyFormer (Ma et al., 2024) extract various frequency information via polynomial approximation like polynomial GNNs. Specformer (Bo et al., 2023) develops learnable filter bases, offering greater expressiveness compared to polynomials with fixed bases. However, such spectral filters still struggle to achieve the desired frequency response due to their limited focus on the specific first-order spectrum.
@@ 3.2 Graph Filter
Figure 2: Overview of GrokFormer. In addition to the use of self-attention to capture global information in the spatial domain, a novel Graph Fourier KAN is proposed in GrokFormer the achieve global graph modeling in the spectral domain. This design enables a strong adaptability in both polynomial order and graph spectrum, offering superior expressive power in capturing diverse graph frequency signals. GrokFormer synthesizes the spatial and spectral representations by a standard summation and normalization layer, followed by a Feed-Forward Network (FFN) layer for prediction.
@@ 3.4 Kolmogorov-Arnold Network
KAN is grounded in the Kolmogorov-Arnold representation theorem (Kolmogorov, 1957), which states that for a function $f$:
@@ 4 Methodology
GrokFormer is a novel GT framework empowered by a Graph Fourier Kolmogorov-Arnold Network (KAN)-based spectral graph convolutional filter, as shown in Figure 2. Graph Fourier KAN in GrokFormer is devised in a way that can adaptively learn diverse frequency signals from a wide range of polynomial order and graph spectrum, going beyond the self-attention mechanism in GTs.
@@ 4.1 The Proposed GrokFormer Filter
Theoretical Analysis. Below we show theoretically that our proposed filter can flexibly fit graph patterns of any order $K$ and graph spectrum.
@@
Our graph filter $h(\lambda)$ is learnable in both polynomial order and graph spectrum:
@@
where the polynomial order $k$ is adaptively determined by coefficient $\alpha_{k}$ while the spectrum $\lambda$ at the specific order $k$ is adaptively determined by coefficients $a_{km}$ and $b_{km}$.
@@ 4.2 Network Architecture of GrokFormer
Homophilic Datasets Heterophilic Datasets
Cora Citeseer Pubmed Photo Physics Penn94 Chameleon Squirrel Actor Texas
Spatial-based GNNs
GCN 87.14$\pm$1.01 79.86$\pm$0.67 86.74$\pm$0.27 88.26$\pm$0.73 97.74$\pm$0.35 82.47$\pm$0.27 59.61$\pm$2.21 46.78$\pm$0.87 33.23$\pm$1.16 77.38$\pm$3.28
GAT 88.03$\pm$0.79 80.52$\pm$0.71 87.04$\pm$0.24 92.94$\pm$0.44 97.82$\pm$0.28 81.53$\pm$0.55 63.13$\pm$1.93 44.49$\pm$0.88 33.93$\pm$2.47 80.82$\pm$2.13
H2GCN 87.96$\pm$0.37 80.90$\pm$1.21 89.18$\pm$0.28 95.45$\pm$0.67 97.19$\pm$0.13 81.31$\pm$0.60 61.20$\pm$4.28 39.53$\pm$0.88 36.31$\pm$2.58 91.89$\pm$3.93
HopGNN 88.68$\pm$1.06 80.38$\pm$0.68 89.15$\pm$0.35 94.49$\pm$0.33 97.86$\pm$0.16 OOM 65.25$\pm$3.49 57.83$\pm$2.11 39.33$\pm$2.79 89.15$\pm$4.04
Spectral-based GNNs
ChebyNet 86.67$\pm$0.82 79.11$\pm$0.75 87.95$\pm$0.28 93.77$\pm$0.32 97.25$\pm$0.78 81.55$\pm$0.26 59.28$\pm$1.25 40.55$\pm$0.42 37.61$\pm$0.89 86.22$\pm$2.45
GPRGNN 88.57$\pm$0.69 80.12$\pm$0.83 88.46$\pm$0.33 93.85$\pm$0.28 97.25$\pm$0.13 81.38$\pm$0.16 67.28$\pm$1.09 50.15$\pm$1.92 39.92$\pm$0.67 92.95$\pm$1.31
BernNet 88.52$\pm$0.95 80.09$\pm$0.79 88.48$\pm$0.41 93.63$\pm$0.35 97.36$\pm$0.30 82.47$\pm$0.21 68.29$\pm$1.58 51.35$\pm$0.73 41.79$\pm$1.01 93.12$\pm$0.65
JacobiConv 88.98$\pm$0.46 80.78$\pm$0.79 89.62$\pm$0.41 95.43$\pm$0.23 97.56$\pm$0.28 83.35$\pm$0.11 74.20$\pm$1.03 57.38$\pm$1.25 41.17$\pm$0.64 93.44$\pm$2.13
HiGCN 89.23$\pm$0.23 81.12$\pm$0.28 89.95$\pm$0.13 95.33$\pm$0.37 97.65$\pm$0.35 OOM 68.47$\pm$0.45 51.86$\pm$0.42 41.81$\pm$0.52 92.15$\pm$0.73
Graph Transformers
Transformer 76.29$\pm$1.95 76.68$\pm$1.21 86.66$\pm$0.50 89.58$\pm$1.05 OOM OOM 45.21$\pm$2.01 33.17$\pm$1.32 39.95$\pm$.64 88.75$\pm$6.30
NodeFormer 87.32$\pm$0.92 79.56$\pm$1.10 89.24$\pm$0.23 95.27$\pm$0.22 96.45$\pm$0.28 69.66$\pm$0.83 56.34$\pm$1.11 43.42$\pm$1.62 34.62$\pm$1.82 84.63$\pm$3.47
SGFormer 87.87$\pm$2.67 79.62$\pm$1.63 89.07$\pm$0.14 94.34$\pm$0.23 97.96$\pm$0.81 76.65$\pm$0.49 61.44$\pm$1.37 45.82$\pm$2.17 41.69$\pm$0.63 92.46$\pm$1.48
NAGphormer 88.15$\pm$1.35 80.12$\pm$1.24 89.70$\pm$0.19 95.49$\pm$0.11 97.85$\pm$0.26 73.98$\pm$0.53 54.92$\pm$1.11 48.55$\pm$2.56 40.08$\pm$1.50 91.80$\pm$1.85
Specformer 88.57$\pm$1.01 81.49$\pm$0.94 90.61$\pm$0.23 95.48$\pm$0.32 97.75$\pm$0.53 84.32$\pm$0.32 74.72$\pm$1.29 64.64$\pm$0.81 41.93$\pm$1.04 88.23$\pm$0.38
PolyFormer(Cheb) 87.67$\pm$1.28 81.80$\pm$0.76 90.68$\pm$0.31 94.08$\pm$1.37 98.08$\pm$0.27 79.27$\pm$0.26 60.17$\pm$1.39 44.98$\pm$3.03 41.51$\pm$0.71 89.02$\pm$5.44
GrokFormer 89.57$\pm$1.43 81.92$\pm$1.25 91.39$\pm$0.51 95.52$\pm$0.52 98.31$\pm$0.18 83.59$\pm$0.26 75.58$\pm$1.73 65.12$\pm$1.59 42.98$\pm$1.48 94.59$\pm$2.08
@@ 4.3 Complexity and Scalability Analysis
Complexity. Firstly, like previous methods (Bo et al., 2023, 2024), GrokFormer also needs spectral decomposition, which is done offline in the preprocessing step and has the complexity of $O(N^{3})$. Secondly, GrokFormer’s forward process involves an embedding layer with the complexity of $O(Nd)$, efficient self-attention with complexity of $O(Nd^{2})$, filter base learning with complexity of $O(KNM)$, and graph convolution with complexity of $O(N^{2}d)$, resulting in an overall forward complexity of $O(Nd(N+d)+KNM)$.
Scalability. In large graphs, GrokFormer can use Sparse Generalized Eigenvalue (SGE) algorithms, as outlined in earlier studies (Cai et al., 2021; Bo et al., 2023, 2024), to compute $q$ eigenvalues and corresponding eigenvectors, in which case the decomposition complexity and the forward complexity will reduce to $O(N^{2}q)$ ($q\ll N$) and $O(Nd^{2}+KqM+Nqd)$, respectively. Empirical results for computational cost can be found in Appendix B.4.

### v3

@@ Abstract
Graph Transformers (GTs) have demonstrated remarkable performance in graph representation learning over popular graph neural networks (GNNs). However, self–attention, the core module of GTs, preserves only low-frequency signals in graph features, leading to ineffectiveness in capturing other important signals like high-frequency ones. Some recent GT models help alleviate this issue, but their flexibility and expressiveness are still limited since the filters they learn are fixed on predefined graph spectrum or spectral order. To tackle this challenge, we propose a Graph Fourier Kolmogorov-Arnold Transformer (GrokFormer), a novel GT model that learns highly expressive spectral filters with adaptive graph spectrum and spectral order through a Fourier series modeling over learnable activation functions. We demonstrate theoretically and empirically that the proposed GrokFormer filter offers better expressiveness than other spectral methods. Comprehensive experiments on 11 real-world node classification datasets across various domains, scales, and graph properties, as well as 5 graph classification datasets, show that GrokFormer outperforms state-of-the-art GTs and GNNs. Our code is available at https://github.com/GGA23/GrokFormer.
Graph Transformers, Kolmogorov-Arnold Networks, Fourier Series, Spectral Graph Neural Networks
@@ 1 Introduction
Graph neural networks (GNNs), which jointly encode graph structures and node features, have been emerging as an effective generic tool for graph-structured learning problems (Yi et al., 2023; Qiao et al., 2024). Despite their effectiveness, popular GNNs are often limited by issues like over-smoothing (Oono & Suzuki, 2020) and over-squashing (Topping et al., 2022). On the other hand, graph Transformers (GTs) use a Transformer-based architecture (Vaswani et al., 2017) to learn graph representations. Due to its strong capability in capturing long-range dependencies among graph nodes, it offers a potential solution to address the issues in popular GNNs.
@@
To address this issue, inspired by polynomial GNNs (e.g., ChebyNet (Defferrard et al., 2016), GPRGNN(Chien et al., 2021), BernNet(He et al., 2021),JacobiConv (Wang & Zhang, 2022)) and A2GCN (Ai et al., 2024), some recent methods are dedicated to capturing various frequency signals by order-$K$ polynomial approximation. For example, FeTA (Bastos et al., 2022) and PolyFormer (Ma et al., 2024) learn the coefficients for order-$K$ polynomial bases (e.g., Chebyshev, Monomial, or Bernstein basis) through the self-attention mechanism. However, these polynomial filters are typically approximated via $K$ predefined bases with specific frequency responses as illustrated in Figure LABEL:fig:movtivation(a), and thus, they have a receptive field of size $K$ in information passing, leading to a locality modeling and limited flexibility and expressivity. Consequently, they are often unable to fit complicated graph filters well, as shown by the failure of the prevalent Bernstein polynomial in fitting a low-comb filter in Figure LABEL:fig:movtivation(b). To effectively encode such spectrum information and achieve a more global graph modeling, Specformer (Bo et al., 2023) performs self-attention over the $N$ eigenvalues after positional encoding to build learnable filter bases. As shown in Figure LABEL:fig:movtivation(a), the frequency response of its filter bases can be arbitrary on the first-order graph Laplacian spectrum. Although it shows impressive effectiveness, its filter learning has a computational complexity of $O(N^{2})$, making it difficult to simultaneously capture higher-order spectral information, and thus misses some important frequency components embedded in higher-order spectrum. Therefore, the learning capacity of the Specformer filter is limited to the specific first-order spectrum and thus it struggles to fit the complicated low-comb filter in Figure LABEL:fig:movtivation(b). Accordingly, the key question we ask here is: can we have a GT that efficiently and flexibly extracts rich frequency signals across the multi-order spectrum of the graph Laplacian?
@@
To answer this question, we propose a novel GT model, called Graph Fourier Kolmogorov-Arnold Transformers (GrokFormer), which provides an efficient approach for learning order- and spectrum-adaptive graph filter for GTs. In particular, motivated by Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024), GrokFormer leverages learnable activation functions modeled as Fourier series over $K$-order graph Laplacian spectrum, producing $K$ adaptive filter bases. As shown in Figure LABEL:fig:movtivation(a), these bases can capture any frequency response across the spectrum of both low and high orders (i.e., from 1st to $K$th orders). Furthermore, we devise learnable order coefficients to assign varying importance to the $K$ filter bases, enabling an adaptive adjustment in fitting the graph spectral order. The resulting filter in GrokFormer is adaptive/learnable in both graph spectrum and spectral order, having better adaptivity than existing popular filters, as shown in Table 1. In doing so, GrokFormer filter can more flexibly capture a broader range frequency responses, offering significantly better expressiveness than other filters (see Figure LABEL:fig:movtivation(b)). The main contributions are as follows:
@@
  • We propose a novel GT model, GrokFormer, that can effectively capture a wide range of frequency signals in an order- and spectrum-adaptive manner. To the best of our knowledge, this is the first GT model that has a learnable filter in both graph spectrum and spectral order.
@@ 2 Related Work
Graph Neural Networks. Existing GNNs are mainly divided into two main streams: spatial-based and spectral-based methods. Spatial-based GNNs, like GCN (Kipf & Welling, 2017), SGC (Wu et al., 2019) and GAT (Veličković et al., 2018), update node representations by aggregating information from neighbors. By stacking multiple layers, they may learn long-range dependencies but suffer from over-smoothing and over-squashing. Some improved spatial methods, such as H2GCN (Zhu et al., 2020), HopGNN (Chen et al., 2023c) and SHGCN (Yan et al., 2025) propose to combine first-hop and multi-hop neighborhood representations. Other studies (Xu et al., 2019; Dong et al., 2021) point out from a spectral perspective that GCN only considers the first-order Chebyshev polynomial, which acts as a low-pass filter. Subsequently, various spectral-based GNNs have been proposed, such as, GPRGNN (Chien et al., 2021), BernNet (He et al., 2021) and JacobiConv (Wang & Zhang, 2022) learn arbitrary graph spectral filters by order-$K$ polynomial approximation. HiGCN (Huang et al., 2024) uses Flower-Petals Laplacians in simplicial complexes to learn polynomial filters across varying topological scales. However, the information passing in these polynomial models is local, and their filters with fixed bases have limited learning ability.
Graph Transformers. Compared to GNNs, the attention weights in Transformers can be viewed as a weighted adjacency matrix of a fully connected graph, capturing long-range dependencies. Some GTs combine both and are popular in graph representation learning, such as Graphormer (Ying et al., 2021), GraphGPS (Rampášek et al., 2022), GRIT (Ma et al., 2023), SAT (Chen et al., 2022), NodeFormer (Wu et al., 2022), NAGphormer (Chen et al., 2023a), GCFormer (Chen et al., 2024b), and SGFormer (Wu et al., 2024) are proposed by incorporating various graph structural information into the Transformer architecture. However, these GTs are limited by the inherent low-pass nature of the self-attention mechanism (Bastos et al., 2022). Advanced GTs have increasingly focused on capturing various frequency signals to tackle the issue. SignGT (Chen et al., 2023b) designs a signed self-attention mechanism to capture low- and high-frequency signals. FeTA (Bastos et al., 2022) and PolyFormer (Ma et al., 2024) extract various frequency information via polynomial approximation like polynomial GNNs. Specformer (Bo et al., 2023) develops learnable filter bases, offering greater spectral expressiveness compared to polynomials with fixed bases. However, such spectral filters still struggle to achieve the desired frequency response due to their limited focus on the specific first-order spectrum.
@@ 3.2 Graph Filter
Figure 2: Overview of GrokFormer. In addition to the use of self-attention to capture global information in the spatial domain, a novel Graph Fourier KAN is proposed in GrokFormer the achieve global graph modeling in the spectral domain. This design enables a strong adaptability in both spectral order and graph spectrum, offering superior expressive power in capturing diverse graph frequency signals. GrokFormer synthesizes the spatial and spectral representations by a standard summation and normalization layer, followed by a Feed-Forward Network (FFN) layer for prediction.
@@ 3.4 Kolmogorov-Arnold Network
KAN is grounded in the Kolmogorov-Arnold representation theorem (Kolmogorov, 1957; Ismayilova & Ismailov, 2024), which states that for a function $f$:
@@ 4 Methodology
GrokFormer is a novel GT framework empowered by a Graph Fourier Kolmogorov-Arnold Network (KAN)-based spectral graph convolutional filter, as shown in Figure 2. Graph Fourier KAN in GrokFormer is devised in a way that can adaptively learn diverse frequency signals from a wide range of spectral order and graph spectrum, going beyond the self-attention mechanism in GTs.
@@ 4.1 The Proposed GrokFormer Filter
Theoretical Analysis. Below we show theoretically that our proposed filter can flexibly fit graph patterns of any spectral order $K$ and graph spectrum.
@@
Our graph filter $h(\lambda)$ is learnable in both spectral order and graph spectrum:
@@
where the spectral order $k$ is adaptively determined by coefficient $\alpha_{k}$ while the spectrum $\lambda$ at the specific order $k$ is adaptively determined by coefficients $a_{km}$ and $b_{km}$.
@@ 4.3 Complexity and Scalability Analysis
Complexity. Firstly, like previous methods (Bo et al., 2023, 2024), GrokFormer also needs spectral decomposition, which is done offline in the preprocessing step and has the complexity of $O(N^{3})$. Secondly, GrokFormer’s forward process involves an embedding layer with the complexity of $O(Nd^{2})$, efficient self-attention with complexity of $O(Nd^{2})$, filter base learning with complexity of $O(KNM)$, and graph convolution with complexity of $O(N^{2}d)$. Note that explicitly constructing the spectral filter matrix $\mathbf{U}diag(h(\lambda))\mathbf{U}^{\top}$ in Eq. (8) incurs a high computational cost of $O(N^{3})$. To address this, we leverage matrix associativity and compute $\mathbf{U}^{\top}\mathbf{X}$ first, which reduce the complexity to $O(N^{2}d)$. As a result, the overall forward pass complexity of GrokFormer is $O(Nd(N+2d)+KNM)$.
Scalability. In large graphs, GrokFormer can use Sparse Generalized Eigenvalue (SGE) algorithms, as outlined in earlier studies (Cai et al., 2021; Bo et al., 2023, 2024), to compute $q$ eigenvalues and corresponding eigenvectors, in which case the decomposition complexity and the forward complexity will reduce to $O(N^{2}q)$ ($q\ll N$) and $O(2Nd^{2}+KqM+Nqd)$, respectively. Empirical results for computational cost can be found in Section 5.5.
Homophilic Datasets Heterophilic Datasets
Cora Citeseer Pubmed Photo WikiCS Physics Penn94 Chameleon Squirrel Actor Texas
Spatial-based GNNs
GCN 87.14$\pm$1.01 79.86$\pm$0.67 86.74$\pm$0.27 88.26$\pm$0.73 82.32$\pm$0.69 97.74$\pm$0.35 82.47$\pm$0.27 59.61$\pm$2.21 46.78$\pm$0.87 33.23$\pm$1.16 77.38$\pm$3.28
GAT 88.03$\pm$0.79 80.52$\pm$0.71 87.04$\pm$0.24 90.94$\pm$0.68 83.22$\pm$0.78 97.82$\pm$0.28 81.53$\pm$0.55 63.13$\pm$1.93 44.49$\pm$0.88 33.93$\pm$2.47 80.82$\pm$2.13
H2GCN 87.96$\pm$0.37 80.90$\pm$1.21 89.18$\pm$0.28 95.45$\pm$0.67 83.45$\pm$0.26 97.19$\pm$0.13 81.31$\pm$0.60 61.20$\pm$4.28 39.53$\pm$0.88 36.31$\pm$2.58 91.89$\pm$3.93
HopGNN 88.68$\pm$1.06 80.38$\pm$0.68 89.15$\pm$0.35 94.49$\pm$0.33 84.73$\pm$0.59 97.86$\pm$0.16 OOM 65.25$\pm$3.49 57.83$\pm$2.11 39.33$\pm$2.79 89.15$\pm$4.04
Spectral-based GNNs
ChebyNet 86.67$\pm$0.82 79.11$\pm$0.75 87.95$\pm$0.28 93.77$\pm$0.32 82.95$\pm$0.45 97.25$\pm$0.78 81.09$\pm$0.33 59.28$\pm$1.25 40.55$\pm$0.42 37.61$\pm$0.89 86.22$\pm$2.45
GPRGNN 88.57$\pm$0.69 80.12$\pm$0.83 88.46$\pm$0.33 93.85$\pm$0.28 82.58$\pm$0.89 97.25$\pm$0.13 81.38$\pm$0.16 67.28$\pm$1.09 50.15$\pm$1.92 39.92$\pm$0.67 92.95$\pm$1.31
BernNet 88.52$\pm$0.95 80.09$\pm$0.79 88.48$\pm$0.41 93.63$\pm$0.35 83.56$\pm$0.61 97.36$\pm$0.30 82.47$\pm$0.21 68.29$\pm$1.58 51.35$\pm$0.73 41.79$\pm$1.01 93.12$\pm$0.65
JacobiConv 88.98$\pm$0.46 80.78$\pm$0.79 89.62$\pm$0.41 95.43$\pm$0.23 84.13$\pm$0.49 97.56$\pm$0.28 83.35$\pm$0.11 74.20$\pm$1.03 57.38$\pm$1.25 41.17$\pm$0.64 93.44$\pm$2.13
HiGCN 89.23$\pm$0.23 81.12$\pm$0.28 89.95$\pm$0.13 95.33$\pm$0.37 83.14$\pm$0.78 97.65$\pm$0.35 OOM 68.47$\pm$0.45 51.86$\pm$0.42 41.81$\pm$0.52 92.15$\pm$0.73
Graph Transformers
Transformer 71.83$\pm$1.68 70.55$\pm$1.20 86.66$\pm$0.50 89.58$\pm$1.05 77.36$\pm$1.25 OOM OOM 45.21$\pm$2.01 33.17$\pm$1.32 39.95$\pm$0.64 88.75$\pm$6.30
GraphGPS 83.42$\pm$1.22 75.87$\pm$0.71 86.62$\pm$0.53 94.35$\pm$0.25 79.26$\pm$0.57 97.60$\pm$0.05 OOM 46.07$\pm$1.51 34.14$\pm$0.73 37.68$\pm$0.94 83.71$\pm$5.85
NodeFormer 87.32$\pm$0.92 79.56$\pm$1.10 89.24$\pm$0.23 95.27$\pm$0.22 81.03$\pm$0.94 96.45$\pm$0.28 69.66$\pm$0.83 56.34$\pm$1.11 43.42$\pm$1.62 34.62$\pm$1.82 84.63$\pm$3.47
SGFormer 87.87$\pm$2.67 79.62$\pm$1.63 89.07$\pm$0.14 94.34$\pm$0.23 82.71$\pm$0.56 97.96$\pm$0.81 76.65$\pm$0.49 61.44$\pm$1.37 45.82$\pm$2.17 41.69$\pm$0.63 92.46$\pm$1.48
NAGphormer 88.15$\pm$1.35 80.12$\pm$1.24 89.70$\pm$0.19 95.49$\pm$0.11 83.41$\pm$0.34 97.85$\pm$0.26 73.98$\pm$0.53 54.92$\pm$1.11 48.55$\pm$2.56 40.08$\pm$1.50 91.80$\pm$1.85
Specformer 88.57$\pm$1.01 81.49$\pm$0.94 90.61$\pm$0.23 95.48$\pm$0.32 85.15$\pm$0.63 97.75$\pm$0.53 84.32$\pm$0.32 74.72$\pm$1.29 64.64$\pm$0.81 41.93$\pm$1.04 88.23$\pm$0.38
PolyFormer 87.67$\pm$1.28 81.80$\pm$0.76 90.68$\pm$0.31 94.08$\pm$1.37 83.62$\pm$0.17 98.08$\pm$0.27 79.27$\pm$0.26 60.17$\pm$1.39 44.98$\pm$3.03 41.51$\pm$0.71 89.02$\pm$5.44
GrokFormer 89.57$\pm$1.43 81.92$\pm$1.25 91.39$\pm$0.51 95.52$\pm$0.52 85.57$\pm$0.65 98.31$\pm$0.18 83.59$\pm$0.26 75.58$\pm$1.73 65.12$\pm$1.59 42.98$\pm$1.48 94.59$\pm$2.08